### v1

@@ V.1.3 Training calculations, emulations, and benchmark calculations
When computing or emulating on-shell $T_{nonBorn}$ amplitude, the $E$ variable in the resolvent operator and the real $E_{rel}$ parameter in the sources are related via $E=E_{rel}-B_{2}$; the potential parameters in the $H$ and sources are undoubtedly the same. However, when performing the training calculations in the form of Eqs. (15) and (16) to collect $|\Psi^{tr}_{\alpha}\rangle$ and $|\tilde{\Psi}^{tr}_{\alpha}\rangle$, the $E$ variable, being complex generally, is treated as independent of $E_{rel}$.
@@
In addition to the elastic scattering amplitude, we also study $\mathcal{A}$ with fixed sources. In these cases, we assign particular values for $E_{rel}$ and the parameters of the $V_{1}$ and $V_{4}$ that are inside the sources (see Eqs. (55) and (57)) and fix them, while varying the $E$ and the $H$’s parameters in the resolvent operator. These calculations resemble computing response functions when $E$ is along the real axis with $\mathrm{Im}\,E=0^{+}$. This type of emulator can also be used to access the compressed spectra and their dependence on the $H$ parameters because the spectrum, as a property of $H$, should be accessible by any sources (or “probes” in terms of response function terminology).
@@
The training solutions are spatially localized and thus can be solved using an integrable basis, such as a Harmonic oscillator basis. However, we work in the momentum space in which the $G_{0}$ is a diagonal operator, and the Faddeev equations can be simplified so that we deal with only one momentum variable, $P$. The nonzero imaginary part of the complex $E$ smooths out the singular functions in solving the training equations, which would be present with real $E$ above the lowest threshold. The difficulty of dealing with moving singularities is mitigated in the training calculations.
@@
These momentum-space equations are derived and simplified step by step in Appendix C.1. Note that in those equations, the parameters in the sources and those in the $H$ are labeled differently, as there are cases in which we need to treat some of them separately. To solve them, we apply the interpolation method Baye (2015) based on the Lagrange function of the Legendre polynomials to discretize the integral equations and obtain a linear equation system, which is solved numerically using a standard linear algebra package. We increase the mesh points to get the desired high accuracy, which could be improved by adapting the careful choice of mesh points and Gaussian quadratures developed in the CE calculations Deltuva and Fonseca (2012).
@@
We emphasize that the training equations in Eqs. (15) and (16) apply for any finite ranged interactions. In our model with separable potentials, however, those equations are simplified to the forms in Appendix C.1, which are then solved to obtain the training solutions.
@@
For benchmarks, we perform two types of calculations: one is with $E$ in the region of training energies, i.e., the test $E$ is in a similar area of the complex plane as the training energies $E^{tr}_{\alpha}$, while the other is on the real energy axis, corresponding to continuum physics of the interests. The conclusion for the first type of benchmark is the same as in the two-body sector: the emulators, as single or multivariate interpolants, have similar accuracy as that of the training calculations.
@@
The important benchmark is the second type. Note the form of our response-like calculation is new, but the scattering $T$-matrix calculation is well established. As checked analytically at the end of Appendix C.2, our equations and formulas provide the same results as the established calculations for on-shell scattering $T$-matrix.
@@
When performing the second type of benchmark calculations, we apply the interpolation method based on the Lagrange function Baye (2015) of the Legendre polynomial to discretize the equations, including Eqs. (96)–(101), as we do in obtaining training solutions. However, the computation becomes significantly different. When $E$ is larger than the lowest threshold, the kernel in the integral equation develops various singularities, and the solution itself has branch points, which require dedicated treatments. Those details can be found in Appendix C.2. With these real-$E$ benchmark calculations setup, we can also extract the information about the bound and resonance states. The procedures are discussed in Appendix C.3.
@@ V.1.4 Numerical values for the Hamitonians
In this work, we fix $\Lambda_{2}$ and $\Lambda_{4}$ in the two and three-body interactions to $\Lambda_{2}=200$ MeV and $\Lambda_{4}=300$ MeV. $\lambda_{4}$ is varied in $[-0.5,0.5]$. We vary $B_{2}\in[2,10]$ MeV when looking at particle-dimer scattering. Note since $\Lambda_{2}$ is fixed, a $B_{2}$ value determines the $\lambda$ value, and vice versa. Associated with the $B_{2}$ range, $\lambda\in[-1.44,-2.15]$. In addition, we vary $\lambda$ in a wider interval, in which the resonances could appear in the three identical boson systems with unbound dimers.
@@ V.2 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$) with fixed $H$ and sources
The emulation is performed only in the complex $E$ plane, while the $H$ and sources are fixed. The source parameters are set as: $E_{rel}^{s}=1$ MeV, $B_{2}^{s}=10$ MeV, and $\lambda_{4}^{s}=-0.2$. They are specifically labeled with the s superscripts to differentiate them from those in the resolvent operator. We then study three cases with distinct $\lambda$ (and the associated $B_{2}$) and $\lambda_{4}$ values for $H$. We will show the compressed spectra and compare them with the “spectra” extracted using the AAA method mentioned in Sec. II.3. The $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ results are also presented.
@@
Figure 15: Emulations in $E$ using our RBM emulator and the AAA method. The $H$ is fixed by setting $B_{2}=10$ MeV and $\lambda_{4}=-0.2$, while the information for the fixed sources is given in the text. The first two panels, sharing the same legend, show the eigenvalues in the complex plane, extracted from the $\mathrm{Im}\,(E_{\alpha}^{tr})=10$ MeV emulators (1st) and the $\mathrm{Im}\,(E_{\alpha}^{tr})=3$ MeV emulators (2nd). Note that in these two panels, the left and right half of the real axis have different scales to accommodate the deep bound state location. The spectra extracted from the same training data sets but using the AAA method (“3A”) are shown for comparisons. The exact results for the bound states are marked as black $\times$, and exact thresholds are indicated by the short vertical black lines. The insets zoom in to the region close to the branch points. The 3rd panel compares various $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ to the exact results on the real energy axis, while their differences, indicating the emulation errors, are shown in the last panel. Both panels share the same legend showing the information about $\mathrm{Im}\,(E^{tr}_{\alpha})$ and $N_{\mathrm{eff}}$.
@@
For Fig. 15, we work with a $H$, whose $B_{2}=10$ MeV and $\lambda_{4}=-0.2$. We trained two emulators with $(\mathrm{Im}\,E^{tr}_{\alpha},N_{b})=(3,48)$ and $(10,30)$ respectively. Similar to the two-body studies, $\mathrm{Re}\,E^{tr}_{\alpha}$ are evenly sampled in the $[-20,50]$ MeV interval. Note that the effective dimension $N_{\mathrm{eff}}$ (see the legend of the 4th panel), i.e., the number of training points needed, is smaller than $N_{b}$ in each case. Furthermore, we performed analytical continuations of the two training data sets of $\mathcal{A}(E^{tr}_{\alpha})$ using the AAA method and collected the corresponding spectra and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ results.
@@
The 1st and 2nd panels, sharing the same legend, show the spectra of the two emulators separately, together with the corresponding AAA results (“3A”). For benchmarking, we mark the exact bound state with black “$\times$” and the exact thresholds with short vertical lines. To accommodate the deep bound state in the plots, the left and right half of the $x$-axis have different scales. The insets in these plots zoom in on the regions near the two branch points. The corresponding $N_{\mathrm{eff}}$ of both emulations and the AAA calculations can be found in the legend of the 4th panel. The AAA $N_{\mathrm{eff}}$ is computed similarly to the emulator $N_{\mathrm{eff}}$ by counting the number of effective poles in the corresponding $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ excluding those extremely close to the branch points and with tiny residues. The 3rd panel compares th emulation results of $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ with the exact ones on the real energy axis below and above the particle-dimer threshold; their differences, i.e., the emulation errors, are plotted in the last panel. Note the two panels share a legend.
@@
Meanwhile, the amplitude, specifically $-\mathrm{Im}\,\mathcal{A}$ on the real energy axis, is also well reproduced by the emulators across the particle-dimer threshold in the 3rd panel.
The comparisons between the two emulation results reinforce our understanding of how $\mathrm{Im}\,E^{tr}_{\alpha}$ impacts the emulator performance. When reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ from $10$ to $3$ MeV, $N_{\mathrm{eff}}$ increases from $11$ to $16$; the density of the DC states increases; and the emulation errors for $-\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ decreases in the challenging region around the two thresholds. Although it is analytically known that below the particle-dimer threshold, $-\mathrm{Im}\,\mathcal{A}$ should be zero, and above that, it starts increasing, the numerical reproduction of this sharp turn behavior is nontrivial. Around that point, the function can not be approximated by Taylor expansions but by rational approximations.
@@
The similarities between the evaluation and the AAA results, including the spectra and amplitudes, are striking. Note that the two are independent calculations; the only information shared is the training data. I.e., no fine-tuning is invoked to achieve such agreement. Therefore, this three-body case further indicates the close connection between the CEE emulation and the near-optimal rational approximation.
Figure 16: Emulations in $E$ using our RBM emulator and the AAA method. The $H$ is fixed by setting $B_{2}=2$ MeV and $\lambda_{4}=0$, while the sources are the same as in Fig. 15. The presentation follows that in Fig. 15, and thus, see the later caption for the explanation of the current plots.
Fig. 16 shows similar results but for a more numerically challenging case: $B_{2}=2$ MeV (i.e., a shallow dimer) and $\lambda_{4}=0$ in $H$, while the sources are the same as used before. There are two three-body bound states; the deeper one has its binding energy $B_{3}$ about 20 MeV, while the other one is loosely bound, with its $B_{3}=2.06$ MeV—only $0.06$ MeV below the particle-dimer threshold. The inset shows the latter bound state more clearly. The emulators and AAA can reproduce these bound states: with $\mathrm{Im}\,E^{tr}=10$, both emulator and AAA give $B_{3}=2.15$, while with $\mathrm{Im}\,E^{tr}=3$, they give $B_{3}=2.08$ MeV. In addition, the emulators and AAA also produce two correct thresholds and discretized BCTs. However, for $\mathrm{Im}\,E^{tr}=10$, the BCT starting at the $E=0$ threshold is severely discretized. Interestingly, the $N_{\mathrm{eff}}$s here have the same dependence on $\mathrm{Im}\,E^{tr}_{\alpha}$ and almost the same values as in Fig. 15.
From the 3rd panel, we see a sharp peak in the exact calculation caused by two competing factors: when moving toward the particle-dimer threshold from its right side, the fast increase of the amplitudes due to the nearby bound state pole and the rapid decrease of the phase space [^19] This feature was found in other studies with shallow bound states (see, e.g., Ref. Hammer and Phillips (2011)).. Although the approximations can reproduce that behavior qualitatively, the errors are significant in that region, relatively speaking, according to the last panel. This is expected based on the low density of the DC states near the thresholds. In the 4th panel, the errors do not peak at the peak location of the $-\mathrm{Im}\,\mathcal{A}$, but rather to its left. Again, the emulation and AAA errors increase towards the threshold but are further amplified by the shallow three-body bound state, as will be encountered in similar cases.
Figure 17: Emulations in $E$ using our RBM emulator and the AAA method. Here, $\lambda=-0.7$ (corresponding to unbound dimer) and $\lambda_{4}=0$, but the sources are the same as in Fig. 15. Also, see the caption of Fig. 15 for explaining the plots.
@@
The general features of the compressed spectra, their dependence on $\mathrm{Im}\,E^{tr}_{\alpha}$, and the similarities between the emulators and the AAA results are the same as discussed in Figs. 15 and 16. The differences in the details include that (1) the two-body interaction is too weak to support a bound dimer, and thus only a three-particle threshold ($E=0$) exists, which is correctly indicated by the single discretized BCT; (2) no three-body bound state exists, but a three-body resonance, as the only physical state, emerges at the location marked by a blue “$+$” in the first two panels. Its exact eigenenergy is $E_{R}=0.29-0.58\,i$ MeV. The resonance is close to the $E=0$ threshold and, at the same time, broad in the sense that the $|\mathrm{Im}\,E_{R}/\mathrm{Re}\,E_{R}|$ ratio is large; and (3) although one of the eigenstates from the emulations and AAA calculations should be identified as the resonance state, there are states in its proximity. This makes resonance identification more challenging than it is for bound states. Again, we rely on the separation between the DC and physical states regarding their distributions in the complex plane and their sensitivity to changing $\mathrm{Im}\,E^{tr}_{\alpha}$. However, in the current case, a high density of the DC states (or the discretized BCT poles) is needed to sharpen these separations, which requires a small $\mathrm{Im}\,E^{tr}_{\alpha}$ value.
@@
The $-\mathrm{Im}\,\mathcal{A}$ (or the response-like function) results on the real axis, as shown in the 3rd panel, can also be used to identify the resonance, similar to how we identify them from the experimental data. However, this procedure faces the same difficulty as just mentioned: the resonance is near the threshold, which calls for precise control of the threshold physics and, thus, again, a high density of the DC states or the discretized BCT poles.
@@
In the 3rd panel, the emulators reproduce the exact results well, including in the region close to the threshold. The errors’ dependence on $\mathrm{Im}\,E^{tr}_{\alpha}$ is similar to what we have seen previously. Moreover, the similarity between the CEE and AAA results is again seen here. The $N_{\mathrm{eff}}$ quantities are also very close to those in Figs. 15 and 16.
@@
Figure 18: Emulations in $E$ using our RBM emulator and the AAA method. Here, $\lambda=-0.7$ (corresponding to unbound dimer) and $\lambda_{4}=-0.2$, but the sources are the same as in Fig. 15. Also, see the caption of Fig. 15 for explaining the plots.
@@
Figure 19: The results from the reanalysis of the same training data sets used in generating Fig. 18. In the new analysis, the BCTs are rotated to a large angle by hand, as discussed in the text. The BCTs seem bent but are on straight lines on a linear scale. No AAA-related results are shown here, as we don’t redefine the BCTs inside the algorithm.
@@
Fig. 18 shows a case with $\lambda=-0.7$ and $\lambda_{4}=-0.2$ in $H$ and the same sources as before. According to the exact calculations, the system has one deep bound state with $B_{3}=691$ MeV and a near-threshold and broad resonance ($E_{R}=-0.15-0.70$ MeV). Interestingly, $\mathrm{Re}\,E_{R}$ is to the left of the $E=0$ threshold.
@@
All the compressed spectra in the first two panels are pretty interesting. They reproduce the bound state and a discretized BCT with an exponentially clustering behavior as in the previous cases and pointing to the correct threshold. However, no resonance state exists in those results. Meanwhile, for $-\mathrm{Im}\,\mathcal{A}$ on the real axis, both emulators and the AAA method perform as well as before, which shows the peak feature but is extremely close to the threshold.

### v2

@@ V.1.3 Training calculations, emulations, and benchmark calculations
When computing or emulating on-shell $T_{\mathrm{nonBorn}}$ amplitude, the $E$ variable in the resolvent operator and the real $E_{\mathrm{rel}}$ parameter in the sources are related via $E=E_{\mathrm{rel}}-B_{2}$; the potential parameters in the $H$ and sources are undoubtedly the same. However, when performing the training calculations in the form of Eqs. (16) and (17) to collect $|\Psi^{\mathrm{tr}}_{\alpha}\rangle$ and $|\tilde{\Psi}^{\mathrm{tr}}_{\alpha}\rangle$, the $E$ variable, being complex generally, is treated as independent of $E_{\mathrm{rel}}$.
@@
In addition to the elastic scattering amplitude, I also study $\mathcal{A}$ with fixed sources. In these cases, I assign particular values for $E_{\mathrm{rel}}$ and the parameters of the $V_{1}$ and $V_{4}$ that are inside the sources (see Eqs. (58) and (60)) and fix them, while varying the $E$ and the $H$’s parameters in the resolvent operator. These calculations resemble computing response functions when $E$ is along the real axis with $\mathrm{Im}\,E=0^{+}$. This type of emulator can also be used to access the compressed spectra and their dependence on the $H$ parameters because the spectrum, as a property of $H$, should be accessible by any sources (or “probes” in terms of response function terminology).
@@
The training solutions are spatially localized and thus can be solved using an integrable basis, such as a Harmonic oscillator basis. However, I work in the momentum space in which the $G_{0}$ is a diagonal operator, and the Faddeev equations can be simplified so that one deals with only one momentum variable, $P$. The nonzero imaginary part of the complex $E$ smooths out the singular functions in solving the training equations, which would be present with real $E$ above the lowest threshold. The difficulty of dealing with moving singularities is mitigated in the training calculations.
@@
These momentum-space equations are derived and simplified step by step in Appendix C.1. Note that in those equations, the parameters in the sources and those in the $H$ are labeled differently, as there are cases in which I need to treat some of them separately. To solve them, I apply the interpolation method [80] based on the Lagrange function of the Legendre polynomials to discretize the integral equations and obtain a linear equation system, which is solved numerically using a standard linear algebra package. I increase the mesh points to get the desired high accuracy, which could be improved by adapting the careful choice of mesh points and Gaussian quadratures developed in the CE calculations [62].
@@
I emphasize that the training equations in Eqs. (16) and (17) apply for any finite ranged interactions. In the model with separable potentials, however, those equations are simplified to the forms in Appendix C.1, which are then solved to obtain the training solutions.
@@
For benchmarks, I perform two types of calculations: one is with $E$ in the region of training energies, i.e., the test $E$ is in a similar area of the complex plane as the training energies $E^{\mathrm{tr}}_{\alpha}$, while the other is on the real energy axis, corresponding to continuum physics of the interests. The conclusion for the first type of benchmark is the same as in the two-body sector: the emulators, as single or multivariate interpolants, have similar accuracy as that of the training calculations.
@@
The important benchmark is the second type. Note the form of my response-like calculation is new, but the scattering $T$-matrix calculation is well established. As checked analytically at the end of Appendix C.2, my equations and formulas provide the same results as the established calculations for on-shell scattering $T$-matrix.
@@
When performing the second type of benchmark calculations, I apply the interpolation method based on the Lagrange function [80] of the Legendre polynomial to discretize the equations, including Eqs. (98)–(103), as I do in obtaining training solutions. However, the computation becomes significantly different. When $E$ is larger than the lowest threshold, the kernel in the integral equation develops various singularities, and the solution itself has branch points, which require dedicated treatments. Those details can be found in Appendix C.2. With these real-$E$ benchmark calculations setup, I can also extract the information about the bound and resonance states. The procedures are discussed in Appendix C.3.
@@ V.1.4 Numerical values for the Hamitonians
In this work, I fix $\Lambda_{2}$ and $\Lambda_{4}$ in the two and three-body interactions to $\Lambda_{2}=200$ MeV and $\Lambda_{4}=300$ MeV. $\lambda_{4}$ is varied in [0.5,0.5][-0.5,0.5]. I vary $B_{2}\in[2,10]$ MeV when looking at particle-dimer scattering. Note since $\Lambda_{2}$ is fixed, a $B_{2}$ value determines the $\lambda$ value, and vice versa. Associated with the $B_{2}$ range, $\lambda$ varies from 1.44-1.44 to 2.15-2.15. In addition, I vary $\lambda\in[-2.15,1]$, a wider range in which the resonances could appear in the three identical boson systems with unbound dimers. Bear in mind that these ranges define the parameter spaces where I will train emulators in the following demonstrations; I also test the extrapolation capability by choosing testing points in larger ranges, if explicitly stated.
@@ V.2 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$) with fixed $H$ and sources
The emulation is performed only in the complex $E$ plane, while the $H$ and sources are fixed. The source parameters are set as: $E_{\mathrm{rel}}^{s}=1$ MeV, $B_{2}^{s}=10$ MeV, and $\lambda_{4}^{s}=-0.2$. They are specifically labeled with the s superscripts to differentiate them from those in the resolvent operator. I then study three cases with distinct $\lambda$ (and the associated $B_{2}$) and $\lambda_{4}$ values for $H$. I will show the compressed spectra and compare them with the “spectra” extracted using the AAA method mentioned in Sec. II.3. The $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ results are also presented.
@@
Figure 12: Emulations in $E$ using the RBM emulator and the AAA method. The $H$ is fixed by setting $B_{2}=10$ MeV and $\lambda_{4}=-0.2$, while the information for the fixed sources is given in the text. The first two panels, sharing the same legend, show the eigenvalues in the complex plane, extracted from the $\mathrm{Im}\,(E_{\alpha}^{\mathrm{tr}})=10$ MeV emulators (1st) and the $\mathrm{Im}\,(E_{\alpha}^{\mathrm{tr}})=3$ MeV emulators (2nd). Note that in these two panels, the left and right half of the real axis have different scales to accommodate the deep bound state location. The spectra extracted from the same training data sets but using the AAA method (“3A”) are shown for comparisons. The exact results for the bound states are marked as black $\times$, and exact thresholds are indicated by the short vertical black lines. The insets zoom in to the region close to the branch points. The 3rd panel compares various $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ to the exact results on the real energy axis, while their differences, indicating the emulation errors, are shown in the last panel. Both panels share the same legend showing the information about $\mathrm{Im}\,(E^{\mathrm{tr}}_{\alpha})$ and $N_{\mathrm{eff}}$.
@@
For Fig. 12, I work with a $H$, whose $B_{2}=10$ MeV and $\lambda_{4}=-0.2$. I trained two emulators with $(\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha},N_{b})=(3,48)$ and $(10,30)$ respectively. Similar to the two-body studies, $\mathrm{Re}\,E^{\mathrm{tr}}_{\alpha}$ are evenly sampled in the [20,50][-20,50] MeV interval. Note that the effective dimension $N_{\mathrm{eff}}$ (see the legend of the 4th panel), i.e., the number of training points needed, is smaller than $N_{b}$ in each case. Furthermore, I performed analytical continuations of the two training data sets of $\mathcal{A}(E^{\mathrm{tr}}_{\alpha})$ using the AAA method and collected the corresponding spectra and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ results.
@@
The 1st and 2nd panels, sharing the same legend, show the spectra of the two emulators separately, together with the corresponding AAA results (“3A”). For benchmarking, the exact bound states are marked with black $\times$” and the exact thresholds with short vertical lines. To accommodate the deep bound state in the plots, the left and right half of the $x$-axis have different scales. The insets in these plots zoom in on the regions near the two branch points. The corresponding $N_{\mathrm{eff}}$ of both emulations and the AAA calculations can be found in the legend of the 4th panel. The AAA $N_{\mathrm{eff}}$ is computed similarly to the emulator $N_{\mathrm{eff}}$ by counting the number of effective poles in the corresponding $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ excluding those extremely close to the branch points and with tiny residues. The 3rd panel compares th emulation results of $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ with the exact ones on the real energy axis below and above the particle-dimer threshold; their differences, i.e., the emulation errors, are plotted in the last panel. Note the two panels share a legend.
@@
Meanwhile, the amplitude, specifically $\mathrm{Im}\,\mathcal{A}$ on the real energy axis, is also well reproduced by the emulators across the particle-dimer threshold in the 3rd panel.
The comparisons between the two emulation results reinforce my understanding of how $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ impacts the emulator performance. When reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ from $10$ to $3$ MeV, $N_{\mathrm{eff}}$ increases from $11$ to $16$; the density of the DC states increases; and the emulation errors for $\mathrm{Im}\,\begin{bmatrix}\mathcal{A}\end{bmatrix}$ decreases in the challenging region around the two thresholds. Although it is analytically known that below the particle-dimer threshold, $\mathrm{Im}\,\mathcal{A}$ should be zero, and above that, it starts increasing, the numerical reproduction of this sharp turn behavior is nontrivial. Around that point, the function can not be approximated by Taylor expansions but by rational approximations.
@@
The similarities between the emulations and the AAA results, including the spectra and amplitudes, are striking. Note that the two are independent calculations; the only information shared is the training data. I.e., no fine-tuning is invoked to achieve such agreement. Therefore, this three-body case further indicates the close connection between the CEE emulation and the near-optimal rational approximation.
Figure 13: Emulations in $E$ using the RBM emulator and the AAA method. The $H$ is fixed by setting $B_{2}=2$ MeV and $\lambda_{4}=0$, while the sources are the same as in Fig. 12. The presentation follows that in Fig. 12, and thus, see the later caption for the explanation of the current plots.
Fig. 13 shows similar results but for a more numerically challenging case: $B_{2}=2$ MeV (i.e., a shallow dimer) and $\lambda_{4}=0$ in $H$, while the sources are the same as used before. There are two three-body bound states; the deeper one has its binding energy $B_{3}$ about 20 MeV, while the other one is loosely bound, with its $B_{3}=2.06$ MeV—only $0.06$ MeV below the particle-dimer threshold. The inset shows the latter bound state more clearly. The emulators and AAA can reproduce these bound states: with $\mathrm{Im}\,E^{\mathrm{tr}}=10$, both emulator and AAA give $B_{3}=2.15$, while with $\mathrm{Im}\,E^{\mathrm{tr}}=3$, they give $B_{3}=2.08$ MeV. In addition, the emulators and AAA also produce two correct thresholds and discretized BCTs. However, for $\mathrm{Im}\,E^{\mathrm{tr}}=10$, the BCT starting at the $E=0$ threshold is severely discretized. Interestingly, the $N_{\mathrm{eff}}$s here have the same dependence on $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ and almost the same values as in Fig. 12.
From the 3rd panel, one can see a sharp peak in the exact calculation caused by two competing factors: when moving toward the particle-dimer threshold from its right side, the fast increase of the amplitudes due to the nearby bound state pole and the rapid decrease of the phase space [^15] This feature was found in other studies with shallow bound states (see, e.g., Ref. [114]).. Although the approximations can reproduce that behavior qualitatively, the errors are significant in that region, relatively speaking, according to the last panel. This is expected based on the low density of the DC states near the thresholds. In the 4th panel, the errors do not peak at the peak location of the $\mathrm{Im}\,\mathcal{A}$, but rather to its left. Again, the emulation and AAA errors increase towards the threshold but are further amplified by the shallow three-body bound state, as will be encountered in similar cases.
Figure 14: Emulations in $E$ using the RBM emulator and the AAA method. Here, $\lambda=-0.7$ (corresponding to unbound dimer) and $\lambda_{4}=0$, but the sources are the same as in Fig. 12. Also, see the caption of Fig. 12 for explaining the plots.
@@
The general features of the compressed spectra, their dependence on $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$, and the similarities between the emulators and the AAA results are the same as discussed in Figs. 12 and 13. The differences in the details include that (1) the two-body interaction is too weak to support a bound dimer, and thus only a three-particle threshold ($E=0$) exists, which is correctly indicated by the single discretized BCT; (2) no three-body bound state exists, but a three-body resonance, as the only physical state, emerges at the location marked by a blue “++” in the first two panels. Its exact eigenenergy is $E_{R}=0.29-0.58\,i$ MeV. The resonance is close to the $E=0$ threshold and, at the same time, broad in the sense that the $|\mathrm{Im}\,E_{R}/\mathrm{Re}\,E_{R}|$ ratio is large; and (3) although one of the eigenstates from the emulations and AAA calculations should be identified as the resonance state, there are states in its proximity. This makes resonance identification more challenging than it is for bound states. Again, one can rely on the separation between the DC and physical states regarding their distributions in the complex plane and their sensitivity to changing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$. However, in the current case, a high density of the DC states (or the discretized BCT poles) is needed to sharpen these separations, which requires a small $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ value.
@@
The $\mathrm{Im}\,\mathcal{A}$ (or the response-like function) results on the real axis, as shown in the 3rd panel, can also be used to identify the resonance, similar to how one identifies them from the experimental data. However, this procedure faces the same difficulty as just mentioned: the resonance is near the threshold, which calls for precise control of the threshold physics and, thus, again, a high density of the DC states or the discretized BCT poles.
@@
In the 3rd panel, the emulators reproduce the exact results well, including in the region close to the threshold. The errors’ dependence on $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is similar to what I have seen previously. Moreover, the similarity between the CEE and AAA results is again seen here. The $N_{\mathrm{eff}}$ quantities are also very close to those in Figs. 12 and 13.
@@
Figure 15: Emulations in $E$ using the RBM emulator and the AAA method. Here, $\lambda=-0.7$ (corresponding to unbound dimer) and $\lambda_{4}=-0.2$, but the sources are the same as in Fig. 12. Also, see the caption of Fig. 12 for explaining the plots.
@@
Figure 16: The results from the reanalysis of the same training data sets used in generating Fig. 15. In the new analysis, the BCTs are rotated to a large angle by hand, as discussed in the text. The BCTs seem bent but are on straight lines on a linear scale. No AAA-related results are shown here, as I can’t redefine the BCTs inside the algorithm.
@@
Fig. 15 shows a case with $\lambda=-0.7$ and $\lambda_{4}=-0.2$ in $H$ and the same sources as before. According to the exact calculations, the system has one deep bound state with $B_{3}=691$ MeV and a near-threshold and broad resonance ($E_{R}=-0.15-0.70$ MeV). Interestingly, the resonance is located to the left of the $E=0$ threshold—since $\mathrm{Re}\,E_{R}<0$.
@@
All the compressed spectra in the first two panels are pretty interesting. They reproduce the bound state and a discretized BCT with an exponentially clustering behavior as in the previous cases and pointing to the correct threshold. However, no resonance state exists in those results. Meanwhile, for $\mathrm{Im}\,\mathcal{A}$ on the real axis, both emulators and the AAA method perform as well as before, which shows the peak feature but is extremely close to the threshold.